# EgoReasoner: Learning Egocentric 4D Reasoning via Task-Adaptive Structured Thinking

### 06 Mar 2026 v1

@@ 4.3 Main Results
Table 1: Quantitative results on the HD-EPIC benchmark across six reasoning tasks. We report multiple-choice accuracy for both open-source baselines and our two-stage approach. Bold indicates best; underline indicates second best. Full table in Supplementary.
Models Avg. 3D Perception Object Motion
@@
VideoLlama 2 [5] 25.5 17.7 18.8 31.0 44.0 11.0 30.5
LongVA [40] 28.1 32.3 26.6 41.2 34.5 10.2 23.5
LLaVA-Video [41] 20.9 16.3 21.8 30.6 20.0 9.8 27.0
Qwen-2.5-VL-3B [32] 27.0 31.0 18.4 37.8 39.0 16.2 19.5
Qwen-2.5-VL-7B [32] 25.7 22.7 22.6 34.0 33.0 20.2 22.0
@@
Unified Training and Task Synergy. We consolidate all tasks into a unified pipeline with optimal reward strategies from separate training. As shown in table˜1, multi-task training promotes cross-task synergy in both stages. These tasks are exceptionally challenging — Qwen2.5-VL-7B (25.7%) does not outperform the 3B model (27.0%), and both trail the human baseline (91.1%) substantially. We thus use the 3B model for primary experiments.
Our unified Stage 2 model achieves 37.2% average accuracy (+11.8 over 7B), with 59.5% on Object Movement Counting (+26.5% over the best baseline) and 50.4% on Object Location, showing that RFT enhances temporal tracking without degrading spatial perception. Stationary Object Localization remains an outlier — its 8–10 minute videos far exceed the 3–5 minute clips of other tasks, making it difficult for RFT to correct reasoning at this temporal scale [2]. We also observe that 3D Perception tasks benefit more from SFT-stage synergy, while Object Motion tasks gain more from RFT, suggesting SFT establishes spatial-semantic foundations that RFT refines for temporal logic.

### 31 Mar 2026 v2

@@ 4.3 Main Results
Table 1: Quantitative results on the HD-EPIC benchmark across six reasoning tasks. We report multiple-choice accuracy for both closed/open-source baselines and our two-stage approach.
Models Size Avg. 3D Perception Object Motion
@@
VideoLlama 2 [5] 7B 25.5 17.7 18.8 31.0 44.0 11.0 30.5
LongVA [41] 7B 28.1 32.3 26.6 41.2 34.5 10.2 23.5
LLaVA-Video [42] 7B 20.9 16.3 21.8 30.6 20.0 9.8 27.0
VideoLlama 3 [39] 7B 26.0 27.7 22.4 35.6 30.5 18.2 21.5
EgoThinker [22] 7B 27.4 28.3 20.0 37.4 30.0 25.0 23.5
VideoChat-R1 [15] 7B 19.2 17.0 19.2 22.8 21.5 13.4 21.5
Qwen-2.5-VL [32] 3B 27.0 31.0 18.4 37.8 39.0 16.2 19.5
Qwen-2.5-VL [32] 7B 25.7 22.7 22.6 34.0 33.0 20.2 22.0
@@
Unified Training and Task Synergy. We consolidate all tasks into a unified pipeline, incorporating the optimal reward strategies identified during separate training. As shown in table˜1, multi-task training promotes cross-task synergy in both stages. These tasks are exceptionally challenging—Qwen2.5-VL-7B (25.7%) does not outperform the 3B model (27.0%), and both trail the human baseline (91.1%) by a wide margin. We therefore adopt the 3B model for our primary experiments.
Our unified Stage 2 model achieves 37.5% average accuracy (+11.8 over the 7B baseline), with notable gains on Object Movement Counting (59.5%, +14.0 over the strongest baseline) and Object Location (50.4%). These results demonstrate that RFT strengthens temporal tracking without degrading spatial perception. Stationary Object Localization remains an outlier—its 8–10 minute videos far exceed the 3–5 minute clips of other tasks, limiting RFT’s ability to correct reasoning at this temporal scale [2]. We also observe that 3D Perception tasks benefit more from SFT-stage synergy, while Object Motion tasks gain primarily from RFT, suggesting that SFT establishes spatial-semantic foundations which RFT then refines for temporal reasoning.